### v1

# Autonomous Exploration and Semantic Updating of Large-Scale Indoor Environments with Mobile Robots

Abstract
We introduce a new robotic system that enables a mobile robot to autonomously explore an unknown environment, build a semantic map of the environment, and subsequently update the semantic map to reflect environment changes, such as location changes of objects. Our system leverages a LiDAR scanner for 2D occupancy grid mapping and an RGB-D camera for object perception. We introduce a semantic map representation that combines a 2D occupancy grid map for geometry, with a topological map for object semantics. This map representation enables us to effectively update the semantics by deleting or adding nodes to the topological map. Our system has been tested on a Fetch robot. The robot can semantically map a $93m\times 90m$ floor and update the semantic map once objects are moved in the environment.  [^1] Project page is available at: https://irvlutd.github.io/SemanticMapping
I Introduction
Autonomous exploration of unknown environments is an important problem in robotics. It has wide applications in robot search and rescue [1], environment surveillance [2] and service robots [3]. Early work on robot exploration focuses on building 2D occupancy grid maps [4] using LiDAR or sonar sensors [5, 6]. Later works consider using 3D octree maps [7] with RGB or RGB-D Simultaneous Localization and Mapping (SLAM) techniques such as ORB-SLAM [8] or KinectFusion [9]. Although these methods can build a geometric representation or map of an unknown environment, the map does not contain semantic information such as objects in the environment.
A common strategy to inject semantics into a geometric map is to leverage object detection or object segmentation using images, and then fuse the detected or segmented objects into the map. This process is commonly known as semantic mapping [10]. Objects can be represented using simple geometric primitives such as boxes or polygons [11, 12], or point clouds from depth cameras [13], or abstract nodes in a graph [14]. By combining robot exploration and semantic mapping, a semantic map of an environment can be built for downstream tasks such as object-goal navigation.
We observed that existing work on robot exploration and semantic mapping typically builds a map once and do not perform updating of the map. Consequently, these methods cannot handle environment changes such as objects being moved around in the environment. Environment changes are difficult for voxel-based or mesh-based mapping methods, since the voxel space or the 3D mesh needs to be updated.
Figure 1: Our system enables a mobile robot to explore a large-scale unknown environment, build a semantic map and update the map for environment changes.
In this work, we introduce a robotic system that is able to autonomously explore a large-scale indoor environment, build a semantic map of the environment online, and subsequently update the semantic map to account for environment changes. To the best of our knowledge, this is the first system that can perform all these tasks jointly on a physical robot. Specifically, our system first enables a mobile robot to explore an unknown environment and build a 2D occupancy grid map of the environment with a LiDAR scanner. Second, given the 2D map, the system plans a trajectory for the robot to traverse the whole environment. Third, the robot follows the planned trajectory and navigates in the environment. Meanwhile, during navigation, we apply open-vocabulary object detection and segmentation in real-time to images from the RGB-D camera on the robot. The detected objects are added to the 2D occupancy grid as nodes in a graph structure. Consequently, a semantic map of the environment can be constructed once the robot finishes the trajectory. Finally, in order to handle environment changes, the robot can visit the environment again and update the semantic map by deleting or adding object nodes online. Fig.  1 illustrates an overview of our system.
We have tested the system on the 4th floor of the Engineering and Computer Science South (ECSS) Building at the University of Texas at Dallas. Our Fetch robot was able to explore the whole floor and build a $93m\times 90m$ 2D occupancy grid map of the floor. In addition, we have shown that the robot can detect chairs, tables, and doors in the environment and build a semantic map for it. To test the effectiveness of the system on updating the semantic map, we intentionally moved some chairs and tables on the floor and let the robot traverse the environment again. We demonstrate that our system is able to update the semantic map to reflect changes in the environment.
II Related Work
II-A Robot Exploration without Semantics
Early works on robot exploration use LiDAR or sonar sensors to build 2D occupancy grid maps [4] of environments. A 2D occupancy grid contains cells that represent free space, obstacles, and unknown space. In 1997, Yamauchi introduced the concept of frontier in robot exploration [5], which built the foundation for frontier-based robot exploration algorithms [15, 6, 16, 17]. Generally, the frontiers on a map represent the boundaries between the unknown space and the free space. They are potential places for a robot to explore. Frontier-based exploration methods can differ in map representations, algorithms for detecting frontiers, and algorithms for selecting frontiers to visit. For example, [17] uses 3D octree maps [7], and the frontiers are 3D voxels in their approach. Most frontier-based exploration methods focus on building a metric map without semantics, such as a 2D occupancy grid map or a 3D octree map. In our method, we fuse semantic information into a 2D occupancy grid map during robot exploration.
II-B Robot Exploration with Semantics
Semantic information about objects in environments is important for robots to perform downstream tasks such as object-goal navigation [18] and off-road navigation [19]. Therefore, researchers incorporate semantics into robot exploration [20, 11, 21, 22]. Some methods simply utilize object segmentation or object detection to detect objects in the environment during exploration and then add semantic information to the map [20, 11]. Some methods consider semantics as another factor for the robot to explore [21, 22]. These methods jointly model geometry and semantics in robot exploration. In our method, we design a two-phase exploration strategy to deal with large-scale environments with large numbers of objects, where the first phase focuses on exploration of geometry to build a 2D occupancy grid map, and the second phase adds semantics into the 2D map.
II-C Semantic Mapping without Robot Exploration
In robot exploration, an exploration algorithm needs to be designed to control the movement of a robot, which also determines the movement of the sensors on the robot such as LiDAR scanners or RGB-D cameras. Therefore, the robot will have the ability to actively explore the environment. On the other hand, there are a large number of methods in the literature that focus on semantic mapping from offline data [23, 13, 24, 25, 26, 27]. These data can be collected by teleoperation of a mobile robot or using a hand-held camera, such as video sequences from a RGB-D camera. Generally speaking, these methods fuse semantic information about objects into a 3D map from an SLAM method. For example, SemanticFusion [13] combines ElasticFusion [28] for dense reconstruction with a neural network for semantic segmentation. DA-RNN [24] combines KinectFusion [9] and a recurrent neural network for video semantic segmentation. Recently, open-vocabulary object detectors such as GroundingDINO [29] have been utilized for semantic mapping. For example, ConceptGraphs [14] and its hierarchical version [30] build 3D scene graphs with open-vocabulary object detectors. Unlike these methods, the map in our method is built via robot exploration. Consequently, a robot can use the map to visit the environment again to refine and update the semantic information in the environment. Very few works have explored updating of the built semantic map except [12]. Our method improves over [12] in mapping larger environments autonomously with much more objects.
Figure 2: Illustration of our system for autonomous exploration, semantic mapping and map update.
III System
Our system is made up of two phases. In the initial phase, the mobile robot explores the environment and builds the occupancy map incrementally. In the subsequent phase, the robot leverages the built occupancy map to accurately localize itself within the environment and proceeds to construct the semantic map. An overview of the system architecture is presented in Fig. 2.
III-A Autonomous Exploration and Map Building
We have developed an environment exploration method based on the popular frontier exploration module [31], using a dynamic search window to efficiently navigate to frontiers in large, real-world environments.
Dynamic Window Frontier Exploration. We modified the this frontier exploration module [31] by dynamically adjusting the robot’s search area between local and global regions [^2] Code available at https://github.com/IRVLUTD/dynamic-window-frontier-exploration. This method identifies frontiers in the current search area and the robot navigates towards them to progressively map the environment. The search window is adjusted dynamically based on the density of the frontiers available around the robot. If the number of frontiers within the local search radius around the robot falls below a threshold, the search expands to a global radius. Setting this parameter to a large value expands the search region to the entire map. This allows the robot to move progressively ahead, reducing redundant coverage of the environment. The exploration continues until the number of frontiers falls below a global threshold or after a fixed time has elapsed, at which point the exploration terminates and saves the built map. Our improved exploration algorithm is integrated with the GMapping SLAM module [32, 33]. The approach is depicted in Fig. 2(a). During this process, the robot exploration trajectory is recorded which is later used to obtain a traversal plan. The saved occupancy map is then utilized to localize the robot during the semantic map construction.
III-B Autonomous Construction of Semantic Map
We utilize the saved occupancy map and the recorded robot exploration path to plan a trajectory to traverse the whole environment to construct the semantic map.
Figure 3: Environment traversal trajectory planning.
III-B1 Environment Traversing
To construct or update a semantic map, the robot needs to visit places in the environment and identify the objects. To optimize this travel, we first sample a subset of $n$ points $P=\{p_{1},p_{2},\dots,p_{n}\}$ along the robot exploration path (Fig. 3(a)), with a minimum of $2m$ separation between each pair of points (Fig. 3(b)). We then construct a graph $G(P,E)$ with the sampled way points as nodes and the Euclidean distance $d$ between them as corresponding edge weights $w_{ij}=d_{ij}=\|p_{i}-p_{j}\|$. This is depicted in Fig. 3(c). Finally, a traversal trajectory $T_{s}$ is obtained from $P$ using a greedy formulation of the Traveling Salesmen Problem (TSP) as explained in Algorithm 1. The generated trajectory is shown in Fig. 3(d).
Algorithm 1 Greedy Algorithm for TSP
1:Input: Sampled waypoints $P=\{p_{1},p_{2},\dots,p_{n}\}$, where $p_{1}$ is the starting waypoint
2:Output: A sequence of waypoints representing the TSP solution trajectory $T_{s}$
3:Initialize trajectory $T_{s}$ $\leftarrow[p_{1}]$
4:Initialize visited $\leftarrow\{p_{1}\}$
5:for $t\leftarrow 1$ to $n-1$ do
6:     Let $p_{\text{last}}\leftarrow$ last waypoint in the $T_{s}$
7:     $p_{\text{next}}\leftarrow\arg\min_{p_{j}\in P\setminus\text{visited}}w(p_{\text{last}},p_{j})$
8:     Append $p_{\text{next}}$ to $T_{s}$
9:     Add $p_{\text{next}}$ to visited
10:end for
11:Append $p_{1}$ to $T_{s}$$\triangleright$ Return to the starting waypoint
12:Return trajectory $T_{s}$
III-B2 Object Detection and Segmentation
During the traversal, object detection and segmentation is performed in real time to construct the semantic map. For real-time object detection, the system leverages GroundingDINO [29] to detect objects in the environment. It takes in the current RGB image from the robot’s camera, identifies objects, and generates labels and bounding boxes for each detection. These bounding boxes, along with the labels, are used as prompts for MobileSAM [34], a faster version implementation of Segment Anything [35], to generate the object segments.
After obtaining segments from the current frame and its corresponding depth image, we compute the 3D point cloud $\mathbf{P}_{\text{camera}}$ for each segment in camera frame by applying camera intrinsic parameters to the depth. These point clouds are then transformed into map frame of reference through the transformation matrices representing camera pose in robot base frame, $\mathbf{T}_{\text{camera}}\in\mathbb{SE}(3)$ and robot pose in map frame $\mathbf{T}_{\text{robot}}\in\mathbb{SE}(3)$:
$\mathbf{P}_{\text{map}}=\mathbf{T}_{\text{robot}}\cdot\mathbf{T}_{\text{camera}}\cdot\mathbf{P}_{\text{camera}}.$ (1)
Finally, the mean positions of the the point clouds are computed using $\mathbf{P}_{\text{mean}}=\frac{1}{m}\sum_{i=1}^{m}\mathbf{P}_{\text{map}}(i)$ for $m$ points in the point cloud as the 3D positions of the corresponding objects in the map frame .
III-B3 Semantic Map Representation and Construction
Representation. Our semantic map is a hybrid representation. At the lower level, we utilize a 2D occupancy grid map $M$ that captures the spatial information of the free space and obstacles in the environment. At the higher level, we construct a topological map $G_{T}(V)$, where $V=\{v_{1},v_{2},\dots,v_{n}\}$ are the nodes representing uniquely detected objects (Fig. 2(d)). Each node $v_{i}$ has the following attributes.
  • $v_{i}.category$: Object category label
  • $v_{i}.position$: Mean position of the object computed in the reference frame of the occupancy map. This attribute connects the topological map to the occupancy map in a hierarchical way.
  • $v_{i}.confidence$: confidence score of the detected object from object detection
This representation facilitates quicker object association and easy update of the nodes in the event of changes in environment.
Construction. To construct the semantic map, we first load the occupancy map $M$ built in Section III-A and localize the robot in it. Then an empty topological map $G_{T}(V)$, $V=\{\}$ is initialized. Additionally, for each object category (e.g., tables, chairs, doors), a category-specific object association distance threshold $\delta_{\{category\}}$ is defined. The robot then starts tracking the traversal trajectory generated in Section III-B1. Simultaneously, object detection and segmentation module is executed to detect objects in the field of view of the robot and compute their positions in the map $M$ frame of reference as illustrated in Fig. 4(a).
As depicted in Fig. 4(b), for every detected object in the robot field of view, its minimum distance to the existing nodes of the corresponding category is calculated. If this distance $<\delta_{\{category\}}$, then that particular object is considered to be already represented on the map and hence not added to $G_{T}$ again. If the minimum distance is greater than the threshold, then the object is added as a new node to $G_{T}$ as seen in Fig. 4(b). The result of this partial construction step is shown in Fig. 4(c). This process is continued until the robot completes the tracking of the traversal trajectory, with new nodes being added to $G_{T}$ as necessary. In this way, the semantic map is fully constructed, combining the geometric structure of the occupancy map $M$ with the layered semantic information of the topological map.
Figure 4: Adding objects to semantic map.
Updating. Once some changes are made in the environment, we proceed to update the semantic map to reflect these changes. This process is similar to the construction phase, but with additional validation steps to handle existing object removal, displacement and new object addition. First, the constructed semantic map is loaded, which acts as reference for the current state of the environment. As the robot tracks the traversal trajectory again, the detected objects in its current field of view are compared against those expected objects to be in the same field of view in the semantic map. If an expected object is no longer detected, it is assumed to have been removed or moved and the corresponding node is deleted from the map as shown in Fig. 5. Similarly, newly detected objects that do not match any existing nodes within distance threshold $\delta_{\{category\}}$ are added as new nodes, representing either new or relocated objects, updating the semantic map in real-time.
Figure 5: Removing objects from semantic map.
IV Experiments
The experiments are carried out in a large indoor environment measuring approximately $100m\times 100m$. It is made up of 18 corridors with a total traversal length of about $800m$. These corridors were mostly furnished with typical office furniture, including doors, tables, and chairs. Hence these 3 classes, i.e., table, door and chair are considered in the open vocabulary detection and segmentation module. We run the experiments on a Fetch mobile robot equipped with a LiDAR sensor and an RGB-D camera. Visual results can be found in Fig. 2.
Exploration and Mapping. To autonomously explore and map the environment, we integrated the mapping, exploration, and navigation modules.The Gmapping ROS package [33] was used to incrementally build a 0.1m/pixel occupancy grid map, using laser scan data from the Fetch robot. Simultaneously, we ran our custom-built Dynamic Window Frontier Exploration module to identify and select best valued frontier in the search window. The Move Base ROS package [36] was then used to navigate the robot to these frontiers. The robot’s maximum velocity was limited to 0.6 m/s with a 0.7m inflation radius to ensure safe navigation. After 150 minutes of exploration, the environment was fully mapped, and 2,250 trajectory points of exploration were recorded at 0.25Hz.
Environment Traversing. From the 2,250 robot poses recorded during the exploration phase, a subset of 130 poses are sampled with a minimum separation 2m distance between every pair of poses. A greedy travelling salesman problem was formulated with these poses and the sequence of sampled points for environment traversal is obtained as described in Sec. III-B1.
Real-time Semantic Map Construction. To enable real-time semantic mapping, the Fetch robot was equipped with a laptop featuring NVIDIA RTX 4090 GPU, running Ubuntu 20.04 and ROS Noetic. The laptop and Fetch robot were connected via Ethernet for subscribing and publishing the data through ROS topics.
Initially, the environment remained unchanged compared to the state during exploration. The system begins with built occupancy map. Localizing itself using AMCL ROS [36] it progressively builds the semantic map. Subsequently, semantic map updates were tested under the following distinct scenarios involving removal, addition, and moving of the objects in the environment.
Online Update I: 10 chairs were removed from a location while 5 chairs at another location have been moved significantly from their original positions.
Online Update II: 10 chairs were removed from a location. 5 chairs and a table have been added at another location and 5 more chairs at another different location. This update was conducted during the night.
IV-A Quantitative Analysis
Construction Update I Update II
Classes GT Map GT Map GT Map
Tables 45 45 45 39 45 37
Chairs 109 84 99 73 109 69
Doors 210 153 210 155 210 156
TABLE I: Semantic map object count across the construction and update stages. GT: Ground Truth, Map: Mapping result. The table compares the ground truth with the actual results across the construction, update I, and update II scenarios.
The robot traverses the environment, updating the semantic map based on the observed changes. The results of the semantic map construction and updates are presented in Table I. We can see that the detection of chairs and doors is relatively more challenging. During update I, the removal of 10 chairs after the construction phase is reflected in the decrease in the number of chairs from 84 to 73. However, in Update II, only 69 chairs were detected, differing from the expected result due to occlusions. Table II presents the breakdown of object detection performance across the three experimental phases. Doors have a relatively low mAP, mainly due to the it’s partial views while robot is traversing in narrow corridors. For chairs, the accuracy decreases during the update stages, indicating the challenges of occlusions.
Table Chair Door mAP
Scenario TP FP FN TP FP FN TP FP FN Table Chair Door
Construction 36 9 4 84 0 16 140 13 68 0.90 0.84 0.67
Update I 36 3 9 72 1 23 140 15 66 0.80 0.76 0.68
Update II 33 4 6 67 2 29 142 14 68 0.85 0.69 0.68
TABLE II: Object detection performance and mAP analysis. TP: True Positives, FP: False Positives, FN: False Negatives, mAP: Mean Average Precision for detected objects across the construction and update stages.
IV-B Qualitative Analysis
Figure 6: Examples of accurate detection
Figure 7: Representation of various challenging scenarios involved in object detection and semantic mapping.
Fig. 6 illustrates several examples of accurate object detection during our experiments for chairs, tables, and doors. In Fig. 7, we discuss representative challenging scenarios in object detection during robot navigation.
Occlusion Challenges. Objects like tables and chairs being in close proximity to each other, often results in occlusion. It leads to objects not being detected or segmented as a single object. This can be seen in Fig. 7(a) where table and chairs are segmented as table. In Fig. 7(c) table, chair, and a portion of the pillar are segmented together as table. In Fig. 7(b), a chair is partially occluded by the pillar. This can lead to incorrect object associations when the chair is detected from another view. In Fig. 7(e), the chairs in the background are completely occluded by the table and are not observed by the robot. Subsequently they are not associated with the semantic map.
Limitations in Constrained Spaces. The presence of constrained spaces like narrow corridors in the environment can hinder the robot’s ability to view the complete structure of objects such as doors as seen in Fig. 7(f). Many such doors are not detected, leading to reduction in the number of object nodes in the semantic map. This is evident in Table I.
Deviation from the Traversal Trajectory. In Fig. 7(d), large bin in the corridor occludes the objects behind it from the robot view, forces the robot to take a detour. This detour results in some objects being seen by the robot. Such scenarios greatly impact the quality of the semantic mapping, reducing the ability to accommodate environment changes.
Impact of Lighting. Varying lighting conditions can cause shadows and reflections. As seen in Fig. 7(g) and Fig. 7(h), these shadows on the floor are mistook for a door and a table, respectively, reducing the overall detection and semantic map accuracy. In Fig. 7(c), the overall environment brightness is reduced during the night, contributing to poor detection and segmentation.
Structural Misidentification. Walls resemble doors when viewed at certain angles, having similar rectangular structure. Some walls also have additional structures like support bars or cable cord covers, resembling large door handles. Hence, these walls are often misidentified as doors. One such case can be seen in Fig. 7(i).
V Conclusion and Future Work
We introduce a robotic system that enables a mobile robot to autonomously explore an unknown environment, recognize objects in the environment, and subsequently build a semantic map of the environment. The system utilizes a hybrid representation of the semantic map that consists of an occupancy grid map for geometry and a topological map for semantics. This representation enables us to update the semantic map efficiently to reflect the changes in the environment. Experiments conducted in a large indoor environment demonstrate the effectiveness of our system in autonomous exploration and semantic mapping.
Object perception during robot navigation still remains a challenge. Our future work will consider interactive object recognition, in which a robot can plan its actions to detect objects such as looking at some objects, moving towards an object to confirm its existence, and moving around to avoid occlusions.
Acknowledgement
This work was supported by the DARPA Perceptually-enabled Task Guidance (PTG) Program under contract number HR00112220005 and the Sony Research Award Program. The work of T. Summers was supported by the United States Air Force Office of Scientific Research under Grant FA9550-23-1-0424 and the National Science Foundation under Grant ECCS-2047040. We would like to thank our colleague, Jishnu Jaykumar P, for his assistance during the experiments.