### v2

# Autonomous Exploration and Semantic Updating of Large-Scale Indoor Environments with Mobile Robots

Abstract
We introduce a new robotic system that enables a mobile robot to autonomously explore an unknown environment, build a semantic map of the environment, and subsequently update the semantic map to reflect environment changes, such as location changes of objects. Our system leverages a LiDAR scanner for 2D occupancy grid mapping and an RGB-D camera for object perception. We introduce a semantic map representation that combines a 2D occupancy grid map for geometry with a topological map for object semantics. This map representation enables us to effectively update the semantics by deleting or adding nodes to the topological map. Our system has been tested on a Fetch robot, semantically mapping a $93m\times 90m$ and a $9m\times 13m$ indoor environment and updating their semantic maps once objects are moved in the environments.  [^1] Project page with code is available at: https://irvlutd.github.io/SemanticMapping
I Introduction
Autonomous exploration of unknown environments is an important problem in robotics. It has wide applications in robot search and rescue [1], environment surveillance [2] and service robots [3]. Early work on robot exploration focuses on building 2D occupancy grid maps [4] using LiDAR or sonar sensors [5, 6]. Later works consider using 3D octree maps [7] with RGB or RGB-D Simultaneous Localization and Mapping (SLAM) techniques such as ORB-SLAM [8] or KinectFusion [9]. Although these methods can build a geometric representation or map of an unknown environment, the map does not contain semantic information such as objects in the environment.
A common strategy to inject semantics into a geometric map is to leverage object detection or object segmentation using images, and then fuse the detected or segmented objects into the map. This process is commonly known as semantic mapping [10]. Objects can be represented using simple geometric primitives such as boxes or polygons [11, 12], point clouds from depth cameras [13], or abstract nodes on a graph [14]. By combining robot exploration and semantic mapping, a semantic map can be built for downstream tasks such as object-goal navigation [15, 16].
We observed that existing work on robot exploration and semantic mapping typically builds a map once and does not perform updating of the map. Consequently, these methods cannot handle environment changes such as objects moving around in the environment. Environment changes are computationally intensive for voxel-based or mesh-based mapping methods, since the voxel space or the 3D mesh needs to be updated.
Figure 1: Our system enables a mobile robot to explore a large-scale unknown environment, build a semantic map and update the map for environment changes.
In this work, we introduce a robotic system that is able to autonomously explore a large-scale indoor environment, build a semantic map of the environment online, and subsequently update the semantic map to account for environment changes. To our knowledge, this is the first system that can perform all these tasks jointly and fully autonomously on a physical robot. Specifically, our system first enables a mobile robot to explore an unknown environment and build a 2D occupancy grid map of the environment with a LiDAR scanner. Second, given the 2D map, the system plans a trajectory for the robot to revisit the whole environment. Third, the robot follows the planned revisiting trajectory and navigates in the environment. During navigation, we apply an open-vocabulary object detection and an object segmentation model in real-time to images from the RGB-D camera on the robot. The detected objects are added to the topological map as nodes in a graph structure. Consequently, a semantic map of the environment can be constructed once the robot finishes the trajectory. Finally, to handle changes in the environment, the robot can visit the environment again and update the semantic map by removing or adding nodes to the map. Fig.  1 illustrates an overview of our system.
We validated our system in a large-scale $(\sim 8,500m^{2})$ and a medium-scale $(\sim 117m^{2})$ indoor environments, using a Fetch robot. The robot successfully explored, mapped, and updated its semantic map of the scene by detecting and adapting to changes such as moved furniture.
Our main contributions in this work are as follows.
  • End-to-End Autonomy: We integrate unseen environment exploration, mapping, revisiting, and semantic updating into one system that runs autonomously on a physical robot online.
  • Practical Semantic Mapping: Using open-vocabulary object detection and segmentation, we build a hybrid representation of semantic map that efficiently scales to larger spaces.
  • Continuous Environment Adaptation: Our system enables real-time semantic map updates to reflect dynamic changes in the environment.
  • Physical System validation: We deploy and evaluate our system on a physical robot across large and medium-scale indoor environments, demonstrating the above contributions.
II Related Work
II-A Robot Exploration without Semantics
Early works on robot exploration use LiDAR or sonar sensors to build 2D occupancy grid maps [4] of environments. A 2D occupancy grid contains cells that represent free space, obstacles, and unknown space. In 1997, Yamauchi introduced the concept of frontier in robot exploration [5], which built the foundation for frontier-based robot exploration algorithms [17, 6, 18, 19]. Generally, the frontiers on a map represent the boundaries between the unknown space and the free space. They are potential places for a robot to explore. Frontier-based exploration methods can differ in map representations, algorithms for detecting frontiers, and algorithms for selecting frontiers to visit. For example, [19] uses 3D octree maps [7], and the frontiers are 3D voxels in their approach. Most frontier-based exploration methods focus on building a metric map without semantics, such as a 2D occupancy grid map or a 3D octree map. In our method, we fuse semantic information into a 2D occupancy grid map during robot exploration.
II-B Robot Exploration with Semantics
Semantic information about objects in environments is important for robots to perform downstream tasks such as object-goal navigation [15] and off-road navigation [20]. Therefore, researchers incorporate semantics into robot exploration [21, 11, 22, 23]. Some methods simply utilize object segmentation or object detection to detect objects in the environment during exploration and then add semantic information to the map [21, 11]. Some methods consider semantics as another factor for the robot to explore [22, 23]. These methods jointly model geometry and semantics in robot exploration. In our method, we designed a two-phase exploration strategy to deal with large-scale environments with large numbers of objects, where the first phase focuses on exploration of geometry to build a 2D occupancy grid map, and the second phase adds semantics into the 2D map.
II-C Semantic Mapping without Robot Exploration
In robot exploration, an exploration algorithm needs to be designed to control the movement of a robot, which also determines the movement of the sensors on the robot such as LiDAR scanners or RGB-D cameras. Thus, the robot will have the ability to actively explore the environment. There are a large number of methods in the literature that focus on semantic mapping from offline data [24, 13, 25, 26, 27, 28]. This data can be collected by teleoperation of a mobile robot or using a hand-held camera, such as video sequences from a RGB-D camera. Therefore, these systems are not fully autonomous and depend on human control to explore and map the environments.
Generally speaking, these methods fuse semantic information about objects into a 3D map from an SLAM method. For example, SemanticFusion [13] combines ElasticFusion [29] for dense reconstruction with a neural network for semantic segmentation. DA-RNN [25] combines KinectFusion [9] and a recurrent neural network for video semantic segmentation. [12] uses Mask R-CNN [30] for instance segmentation to construct and update a semantic map in real-time using a mobile manipulator system. However, it requires a fully explored environment map as a prerequisite to begin with. Hydra [31] builds and optimizes a 3D scene graph hierarchically from simulated and real-world datasets. Khronos [32] constructs dense spatio-temporal maps that reflect the 3D scene and also account for some dynamic environment changes while testing their method using a physical robot system. However, the robots are controlled manually during this process. Recently, open-vocabulary object detectors such as GroundingDINO [33] have been utilized for semantic mapping. For example, ConceptGraphs [14] and its hierarchical version [34] build 3D scene graphs with open-vocabulary object detectors based on a hand-held scanned dataset of the indoor environment.
Although these works significantly advance in semantic mapping, they all share a common limitation, i.e., they are not fully autonomous. As shown in Table I, prior methods are based on manual control for environmental exploration. Some methods require pre-explored maps, and some do not perform semantic updating of environments. In contrast, our method initially builds the map via robot exploration. Consequently, the robot can use this map to revisit the environment in a planned and greedy manner, while continuously refining and updating the semantic information in the environment. This offers a fully autonomous approach to both mapping and semantic updating.
Method Autonomous Exploration Mapping with Robot Semantic Mapping Semantic Updating
SemanticFusion [13] × × ×
DA-RNN [25] × × ×
Hydra[31] × × ×
ConceptGraphs[14] × × ×
Dengler et al.[12] × ×
Khronos[32] ×
Ours
TABLE I: Comparison of methods for autonomous exploration, semantic mapping and updating (reflecting environment changes).
Figure 2: Illustration of our system for autonomous exploration, semantic mapping and map update.
III System
Our system consists of two phases. In the initial phase, the mobile robot explores the environment and builds the occupancy map incrementally. In the subsequent phase, the robot leverages the built occupancy map to accurately localize itself within the environment and proceeds to construct the semantic map. The motivation to separate the process into two stages, rather than performing simultaneously, is because of computational efficiency and robustness. Attempting to perform both tasks in a single stage could lead to resource constraints and reduced performance, especially for large-scale environments with a large number of objects. Additionally, in large environments, as the number of detected objects increases, the number of map corrections (like loop closures) applied to the detected objects increase, which slows down the mapping process. Evidently, this has also been mentioned in [32]. Therefore, our two-phase approach is more efficient for semantic exploration and mapping of large-scale environments. An overview of the system is presented in Fig. 2.
III-A Autonomous Exploration and Map Building
We have developed an environment exploration method based on the popular frontier exploration module [35], using a dynamic search window to efficiently navigate to frontiers in large real-world environments. When using [35] in large scale environments, as the map size increases, far-away frontiers become insignificant due to high cost, thus leaving a part of the map unexplored. Moreover, tuning the cost parameters in [35] is relatively hard and not intuitive.
Figure 3: Environment traversal trajectory planning.
Dynamic Window Frontier Exploration. We modified the frontier exploration module [35] by dynamically adjusting the robot’s search area between local and global regions [^2] Code available at https://github.com/IRVLUTD/dynamic-window-frontier-exploration. This method identifies frontiers in the current search area and the robot navigates towards them to progressively map the environment. The search window is adjusted dynamically based on the density of the frontiers available around the robot. If the number of frontiers within the local search radius around the robot falls below a threshold, the search expands to a global radius. Setting this parameter to a large value expands the search region to the entire map. This allows the robot to move progressively ahead, reducing redundant coverage of the environment. The exploration continues until the number of frontiers falls below a global threshold or after a fixed time has elapsed, at which point the exploration terminates and saves the built map. Our improved exploration algorithm is integrated with the GMapping SLAM module [36, 37]. The approach is depicted in Fig. 2(a). During this process, the robot exploration trajectory is recorded which is later used to obtain a traversal plan. The saved occupancy map is then utilized to localize the robot during the semantic map construction and update phases.
III-B Autonomous Construction of Semantic Map
We utilize the saved occupancy map and the recorded robot exploration path to plan a trajectory to traverse the whole environment to construct the semantic map.
III-B1 Environment Traversing
To construct or update a semantic map, the robot needs to visit places in the environment and identify the objects. To optimize this travel, we first sample a subset of $n$ points $P=\{p_{1},p_{2},\dots,p_{n}\}$ along the robot exploration path (Fig. 3(a)), with a minimum of $2m$ separation between each pair of points (Fig. 3(b)). We then construct a graph $G(P,E)$ with the sampled way points as nodes and the Euclidean distance $d$ between them as corresponding edge weights $w_{ij}=d_{ij}=\|p_{i}-p_{j}\|$. This is depicted in Fig. 3(c). Finally, a traversal trajectory $T_{s}$ is obtained from $P$ using a greedy formulation of the Traveling Salesmen Problem (TSP) as explained in Algorithm 1. The generated trajectory is shown in Fig. 3(d).
Algorithm 1 Greedy Algorithm for TSP
1:Input: Sampled waypoints $P=\{p_{1},p_{2},\dots,p_{n}\}$, where $p_{1}$ is the starting waypoint
2:Output: A sequence of waypoints representing the TSP solution trajectory $T_{s}$
3:Initialize trajectory $T_{s}$ $\leftarrow[p_{1}]$
4:Initialize visited $\leftarrow\{p_{1}\}$
5:for $t\leftarrow 1$ to $n-1$ do
6:     Let $p_{\text{last}}\leftarrow$ last waypoint in the $T_{s}$
7:     $p_{\text{next}}\leftarrow\arg\min_{p_{j}\in P\setminus\text{visited}}w(p_{\text{last}},p_{j})$
8:     Append $p_{\text{next}}$ to $T_{s}$
9:     Add $p_{\text{next}}$ to visited
10:end for
11:Append $p_{1}$ to $T_{s}$$\triangleright$ Return to the starting waypoint
12:Return trajectory $T_{s}$
III-B2 Object Detection and Segmentation
During the traversal, object detection and segmentation is performed in real time to construct the semantic map. For real-time object detection, the system leverages GroundingDINO [33] to detect objects in the environment. It takes in the current RGB image from the robot’s camera, identifies objects, and generates labels and bounding boxes for each detection. These bounding boxes are used as prompts for MobileSAM [38], a faster version of the Segment Anything Model (SAM) [39], to generate the object segmentation masks.
After obtaining object segments of the current observation and its corresponding depth image, we compute the 3D point cloud $\mathbf{P}_{\text{camera}}$ for each segment in camera frame by applying camera intrinsic parameters to the depth. These points are then transformed to the map frame of reference through the transformation matrices representing camera pose in robot base frame, $\mathbf{T}_{\text{camera}}\in\mathbb{SE}(3)$ and robot pose in map frame $\mathbf{T}_{\text{robot}}\in\mathbb{SE}(3)$:
$\mathbf{P}_{\text{map}}=\mathbf{T}_{\text{robot}}\cdot\mathbf{T}_{\text{camera}}\cdot\mathbf{P}_{\text{camera}}.$ (1)
Finally, the mean position of the point cloud is computed using $\mathbf{P}_{\text{mean}}=\frac{1}{m}\sum_{i=1}^{m}\mathbf{P}_{\text{map}}(i)$ for $m$ points in the point cloud as the 3D position of the corresponding object in the map frame.
III-B3 Semantic Map Representation and Construction
Our semantic map is a hybrid representation. At the lower level, we utilize a 2D occupancy grid map $M$ that captures the spatial information of the free space and obstacles in the environment. At the higher level, we construct a topological map $G_{T}(V)$, where $V=\{v_{1},v_{2},\dots,v_{n}\}$ are the nodes representing uniquely detected objects (Fig. 2(d)). Each node $v_{i}$ has the following attributes.
  • $v_{i}.\text{id}$: A unique id of the object
  • $v_{i}.\text{category}$: Object category label
  • $v_{i}.\text{position}$: Mean position of the object computed in the reference frame of the occupancy map. This attribute connects the topological map to the occupancy map in a hierarchical way.
  • $v_{i}.\text{confidence}$: the confidence score of the detected object from object detection
This representation facilitates fast object association and easy update of the nodes in the event of changes in the environment.
Construction. To construct the semantic map, we first load the occupancy map $M$ built in Section III-A and localize the robot using it. Then an empty topological map $G_{T}(V)$, $V=\{\}$ is initialized. Additionally, for each object category (e.g., tables, chairs, doors), a category-specific object association distance threshold $\delta_{\{\text{category}\}}$ is defined. The robot then starts tracking the traversal trajectory generated in Section III-B1. Simultaneously, object detection and segmentation module is executed to detect objects in the field of view of the robot and compute their positions in the map $M$ frame of reference as illustrated in Fig. 4(a).
As depicted in Fig. 4(b), for every detected object in the robot field of view, its minimum distance to the existing nodes of the corresponding category is calculated. If this distance is small than $\delta_{\{\text{category}\}}$, then that particular object is considered to be already represented on the map and hence not added to the topological map $G_{T}$ again. If the minimum distance is greater than the threshold, then the object is added as a new node to $G_{T}$ as illustrated in Fig. 4(b). The result of this partial construction step is shown in Fig. 4(c). This process is continued until the robot completes the tracking of the traversal trajectory, with new nodes being added to $G_{T}$ after object association. In this way, the semantic map is fully constructed, combining the geometric structure of the occupancy map with the layered semantics of the topological map.
Figure 4: Adding objects to semantic map.
Updating. Once some changes are made to the environment, we update the semantic map to reflect these changes. This process is similar to the construction phase, but with additional validation steps to handle existing object removal, displacement, and new object addition. First, the constructed semantic map is loaded, which acts as a reference for the current state of the environment. As the robot tracks the traversal trajectory again, the detected objects in its current field of view are compared against those objects that are expected to be in the same field of view in the semantic map. If an expected object is no longer detected, it is assumed to have been removed or moved and the corresponding node is deleted from the map as shown in Fig. 5. Similarly, newly detected objects that do not match any existing nodes within the distance threshold $\delta_{\{\text{category}\}}$ are added as new nodes, representing new or relocated objects. This process updates the semantic map in real-time.
Figure 5: Removing objects from semantic map.
IV Experiments
We have conducted experiments in two different environments with different object categories and evaluate the performance of our system in those environments.
Environment A. The fourth floor of the ECSS building at UT Dallas is a large indoor environment measuring approximately $93m\times 90m$. It consists of 18 corridors with a total traversal length of approximately $800m$. These corridors were mostly furnished with typical office furniture, including doors, tables, and chairs. Therefore, we only consider these 3 classes, i.e., table, door and chair for semantic mapping of this environment.
Environment B. This environment is a laboratory environment with a reduced size measuring $9m\times 13m$. This space included a more diverse set of objects such as televisions, trash bins, umbrellas, chairs, cabinets and persons.
Robot. We run the experiments on a Fetch mobile robot equipped with a LiDAR sensor and an RGB-D camera as depicted in Fig. 6.
Figure 6: The Fetch robot maps the environment using a $220^{o}$ laser scanner of $25m$ range, while leveraging depth observations for obstacle avoidance. Its RGB-D camera operates at 30fps, enabling real-time object detection and segmentation for semantic mapping running on the connected GPU laptop.
Exploration and Mapping. To autonomously explore and map the environments, we integrate the mapping, exploration, and navigation modules. The Gmapping ROS package [37] was used to incrementally build a 0.1m/pixel occupancy grid map, using laser scan data from the Fetch robot. Simultaneously, we ran our custom-built Dynamic Window Frontier Exploration module to identify and select best valued frontier in the search window. The move_base ROS package [40] guided the robot to these frontiers by combining a global planner for long term path planning and a local planner that dynamically avoids obstacles not seen during global planning. This ensures safe navigation when any new objects are added or moved around in the environment. The robot’s maximum velocity was limited to 0.6m/s, with a 0.7m inflation radius to prevent it from traveling too close to the obstacles. In Environment A, after 150 minutes of exploration, which also included several stops due to close proximity of persons in narrow corridors, the environment was fully mapped, and 2,250 trajectory points were recorded at 0.25Hz. In Environment B, exploration was completed in approximately 4 minutes, with 212 points recorded at 1Hz.
Environment Traversing. The greedy TSP described in Sect. III-B1 is formulated using the poses recorded during the exploration, resulting in traversal trajectories of 130 poses for Environment A and 15 poses for Environment B. In particular, the time to revisit is significantly reduced compared to the exploration phase. Revisiting took approximately 35 minutes for Environment A and 2 minutes for Environment B. This highlights the efficiency gained through this traversal optimization.
Real-time Construction and Updating of Semantic Maps. To enable real-time semantic mapping, the Fetch robot was equipped with a laptop featuring NVIDIA RTX 4090 GPU, running Ubuntu 20.04 and ROS Noetic. The laptop and Fetch robot were connected via Ethernet for subscribing and publishing the data through ROS topics.
Initially, the environment remained unchanged compared to its state during exploration. The system then begins with the built occupancy map. After localizing itself using AMCL ROS [40], the robot progressively builds the semantic map. Subsequently, semantic map updates were tested by removal, addition, and/or moving the objects in the environments. Visual results of environment A can be found in Fig. 2. More details of it can be found in the supplementary video. Fig. 7 depicts some of the changes in Environment B, reflected in the semantic updating stages.
Figure 7: Visualization of some selected environmental changes in Environment B across two update stages. In Update stage I, 2 trash bins are removed. After that in update stage II, 2 chairs were removed. When robot revisited, these changes are reflected in the updated semantic map.
Env Class Construction Update I Update II
GT Map GT Map GT Map
A Table 45 45 45 39 45* 37
Chair 109 84 99 73 109* 69
Door 170 153 170 155 170 156
B Person 2 2 2* 2 2 2
Trash Bin 2 2 0 0 0 1
Umbrella 1 1 0 0 0 0
Chair 3 3 3 3 1* 1
Television 1 2 1 1 1 1
Cabinet 3 2 3 1 3 2
TABLE II: Semantic map object count across the construction and update stages for A & B Environments. GT: Ground Truth, Map: no. of objects mapped. The table compares the ground truth with the actual results across the construction, update I, and update II scenarios. * indicates that some objects have been relocated from their earlier position.
IV-A Quantitative Analysis
Table II summarizes the number of objects detected in both Environment A and Environment B during the Construction, Update I and Update II phases. In Environment A, we can see that while doors and tables were stably detected, the detection of chairs was relatively more challenging. During update I, the removal of 10 chairs after the construction phase is reflected in the decrease in the number of chairs detected from 84 to 73. However, in Update II, only 69 chairs were detected, with false negatives increasing to 29 as shown in Table III. This was mainly because some relocated chairs were placed in the path of the robot traversal trajectory. As a result, robot took a detour to avoid hitting them, causing it to miss some chairs that were not relocated and previously detected. Moreover, people occupying chairs occluded them, increasing false negatives. In contrast, Environment B has very stable detections, with many classes having 100% precision. However, some categories like cabinets that are dark colored, are challenging to detect in nominal lightning, resulting in them not being detected.
IV-B Qualitative Analysis
Figure 8: Examples of accurate detection and segmentation of chairs, tables and doors in Environment A.
Class Construction Update I Update II
TP FP FN Precision TP FP FN Precision TP FP FN Precision
Environment A
Table 36 9 4 0.9 36 3 9 0.8 33 4 6 0.85
Chair 84 0 16 0.84 72 1 23 0.76 67 2 29 0.69
Door 140 13 28 0.83 140 15 26 0.84 142 14 28 0.84
Environment B
Person 2 0 0 1 2 0 0 1 2 0 0 1
Trash Bin 2 0 0 1 0 0 0 - 0 1 0 0
Umbrella 1 0 0 1 0 0 0 - 0 0 0 -
Chair 3 0 0 1 3 0 0 1 1 0 0 1
Television 1 1 0 0.5 1 0 0 1 1 0 0 1
Cabinet 2 0 1 0.67 1 0 2 0.33 2 0 1 0.67
TABLE III: Semantic map object count across the construction and update stages for A & B Environments. TP: True Positives, FP: False Positives, FN: False Negatives, Precision: Precision for detected objects across the construction and update stages.
In Fig. 9, we show several challenging scenarios in object detection and segmentation, while robot is navigating.
Occlusion Challenges. Close proximity of objects such as tables and chairs causes occlusion and leads to missed detections or incorrect segmentation, as seen in Fig. 9(a). In Fig. 9(c) table, chair, and a portion of the pillar are segmented together as table. The scenario seen in Fig. 9(b), where a chair is partially occluded by the pillar, leads to incorrect object associations when the chair is detected again at another view.
Detection Threshold. Lowering the detection threshold increases false positives, while a higher threshold may miss objects. In Fig. 9(e), chairs in the foreground are undetected at a 0.8 threshold. careful choice of threshold for object class is important for accurate semantic mapping.
Limitations in Constrained Spaces. Narrow corridors limit the robot’s view of objects like doors, as seen in Fig. 9(f). Several of such doors are not detected, leading to a reduction in the number of object nodes in the semantic map. This is evident in Table II.
Figure 9: Representation of various challenging scenarios involved in object detection and semantic mapping.
Deviation from the Traversal Trajectory. In Fig. 9(d), a large bin in the corridor occluded objects, forcing the robot to take a detour. This detour resulted in the robot not seeing some objects. Such scenarios impact the quality of the semantic mapping, reducing the ability to accommodate environmental changes.
Impact of Lighting. Varying lighting conditions can cause shadows and reflections. As seen in Fig. 9(g) and Fig. 9(h), these shadows on the floor were mistook for a door and a table, respectively, reducing the overall detection and semantic map accuracy. In Fig. 9(c), reduced brightness at night degraded detection and segmentation.
Structural Misidentification. Walls resembling doors due to shape or added features (e.g., support bars) often cause misidentification, as seen in Fig. 9(i). This scenario also occurred in Environment B, update II phase, where a cardboard box was identified as a trash bin, leading to a false positive.
V Conclusion and Future Work
We introduce a robotic system that enables a mobile robot to autonomously explore an unknown environment, recognize objects in the environment, and subsequently build a semantic map of the environment. The system utilizes a hybrid representation of the semantic map that consists of an occupancy grid map for geometry and a topological map for semantics. This representation enables us to update the semantic map efficiently to reflect the changes in the environment. Experiments conducted on a mobile manipulatory system in a large- and medium-scale indoor environments demonstrate the effectiveness of our system in autonomous exploration and semantic mapping.
Object perception during robot navigation still remains a challenge. Our future work will consider interactive object recognition, in which a robot can plan its actions to detect objects such as looking at some objects, moving towards an object to confirm its existence for a robust object association, and dynamic planning of traversal to avoid occlusions.
Acknowledgment
This work was supported by the DARPA Perceptually-enabled Task Guidance (PTG) Program under contract number HR00112220005 and the Sony Research Award Program. The work of T. Summers was supported by the United States Air Force Office of Scientific Research under Grant FA9550-23-1-0424 and the National Science Foundation under Grant ECCS-2047040. We would like to thank our colleague, Jishnu Jaykumar P, for his assistance during the experiments.