### v3

# A Modular Robotic System for Autonomous Exploration and Semantic Updating in Large-Scale Indoor Environments

Abstract
We present a modular robotic system for autonomous exploration and semantic updating of large-scale unknown environments. Our approach enables a mobile robot to build, revisit, and update a hybrid semantic map that integrates a 2D occupancy grid for geometry with a topological graph for object semantics. Unlike prior methods that rely on manual teleoperation or precollected datasets, our two-phase approach achieves end-to-end autonomy: first, a modified frontier-based exploration algorithm with dynamic search windows constructs a geometric map; second, using a greedy trajectory planner, environments are revisited, and object semantics are updated using open-vocabulary object detection and segmentation. This modular system, compatible with any metric SLAM framework, supports continuous operation by efficiently updating the semantic graph to reflect short-term and long-term changes such as object relocation, removal, or addition. We validate the approach on a Fetch robot in real-world indoor environments of approximately 8,5008,500m2 and 117117m2, demonstrating robust and scalable semantic mapping and continuous adaptation, marking a fully autonomous integration of exploration, mapping, and semantic updating on a physical robot.  [^1] Project page with code is available at: https://irvlutd.github.io/SemanticMapping
I Introduction
Autonomous exploration of unknown environments is an important problem in robotics. It has wide applications in robot search and rescue [1], environment surveillance [2] and service robots [3]. Early work on robot exploration focuses on building 2D occupancy grid maps [4] using LiDAR or sonar sensors [5, 6]. Later works consider using 3D octree maps [7] with RGB or RGB-D Simultaneous Localization and Mapping (SLAM) techniques such as ORB-SLAM [8] or KinectFusion [9]. Although these methods can build a geometric representation or map of an unknown environment, the map does not contain semantic information such as objects in the environment.
A common strategy to inject semantics into a geometric map is to leverage object detection or object segmentation using images, and then fuse the detected or segmented objects into the map. This process is commonly known as semantic mapping [10]. Objects can be represented using simple geometric primitives such as boxes or polygons [11, 12], point clouds from depth cameras [13], or abstract nodes on a graph [14]. By combining robot exploration and semantic mapping, a semantic map can be built for downstream tasks such as object-goal navigation [15, 16].
We observed that existing work on robot exploration and semantic mapping typically builds a map once and does not perform updating of the map. Consequently, these methods cannot handle environment changes such as objects moving around in the environment. Environment changes are computationally intensive for voxel-based or mesh-based mapping methods, since the voxel space or the 3D mesh needs to be updated.
Figure 1: Our system enables a mobile robot to explore a large-scale unknown environment, build a semantic map and update the map for environment changes.
In this work, we introduce a robotic system that is able to autonomously explore a large-scale indoor environment, build a semantic map of the environment online, and subsequently update the semantic map to account for environment changes. Our system represents a significant step forward in integrating autonomous exploration, long-term semantic mapping, and updating on a physical robot, addressing limitations of prior methods that rely on manual control or posed datasets. Specifically, our system first enables a mobile robot to explore an unknown environment and build a 2D occupancy grid map of the environment with a LiDAR scanner. Second, given the 2D map, the system plans a trajectory for the robot to revisit the whole environment. Third, the robot follows the planned revisiting trajectory and navigates in the environment. During navigation, we apply an open-vocabulary object detection and an object segmentation model in real-time to images from the RGB-D camera on the robot. The detected objects are added to the topological map as nodes in a graph structure. Consequently, a semantic map of the environment can be constructed once the robot finishes the trajectory. Finally, to handle changes in the environment, the robot can visit the environment again and update the semantic map by removing or adding nodes to the map. Fig.  1 illustrates an overview of our system.
We validated our system in a large-scale 8,5008,500m2 and a medium-scale 117117m2 indoor environments, using a Fetch robot. The robot successfully explored, and updated the semantic map by detecting changes such as moved furniture.
Our main contributions in this work are as follows.
  • End-to-End Autonomy: We integrate unseen environment exploration, mapping, revisiting, and semantic updating into a modular system that operates fully autonomously on a physical robot in real-time.
  • Lightweight Object Association: An efficient association strategy that enables fast and scalable semantic map updates.
  • Short and Long-Term Semantic Updating: Our semantic mapping approach captures changes during the current visit as well as changes observed across multiple revisits, enabling the map to reflect both immediate and accumulated environmental updates.
  • Real-World Validation at Scale: We deploy and evaluate our system on a physical robot in large- and medium-scale indoor environments, demonstrating the above contributions.
II Related Work
II-A Robot Exploration without Semantics
Early works on robot exploration use LiDAR or sonar sensors to build 2D occupancy grid maps [4] of environments. A 2D occupancy grid contains cells that represent free space, obstacles, and unknown space. In 1997, Yamauchi introduced the concept of frontier in robot exploration [5], which built the foundation for frontier-based robot exploration algorithms [17, 6, 18, 19]. Generally, the frontiers on a map represent the boundaries between the unknown space and the free space. They are potential places for a robot to explore. Frontier-based exploration methods can differ in map representations, algorithms for detecting frontiers, and algorithms for selecting frontiers to visit. For example, [19] uses 3D octree maps [7], and the frontiers are 3D voxels in their approach. Most frontier-based exploration methods focus on building a metric map without semantics, such as a 2D occupancy grid map or a 3D octree map. In our method, we fuse semantic information into a 2D occupancy grid map during robot exploration.
II-B Robot Exploration with Semantics
Semantic information about objects in environments is important for robots to perform downstream tasks such as object-goal navigation [15] and off-road navigation [20]. Therefore, researchers incorporate semantics into robot exploration [21, 11, 22, 23]. Some methods simply utilize object segmentation or object detection to detect objects in the environment during exploration and then add semantic information to the map [21, 11]. Some methods consider semantics as another factor for the robot to explore [22, 23]. These methods jointly model geometry and semantics in robot exploration. In our method, we designed a two-phase exploration strategy to deal with large-scale environments with large numbers of objects, where the first phase focuses on exploration of geometry to build a 2D occupancy grid map, and the second phase adds semantics into the 2D map.
Method Autonomous Exploration Metric Mapping with Robot Semantic Mapping Semantic Updating
SemanticFusion[13] × × ×
DA-RNN [24] × × ×
Hydra[25] × × ×
ConceptGraphs[14] × × ×
Dengler et al.[12] × ×
Khronos[26] ×
Ours
TABLE I: Comparison of methods for autonomous exploration, metric mapping with a physical robot, semantic mapping and updating (reflecting environment changes).
II-C Semantic Mapping without Robot Exploration
In robot exploration, an exploration algorithm needs to be designed to control the movement of a robot, which also determines the movement of the sensors on the robot such as LiDAR scanners or RGB-D cameras. Thus, the robot will have the ability to actively explore the environment. There are a large number of methods in the literature that focus on semantic mapping of offline data [27, 13, 24, 28, 29, 30]. This data can be collected by teleoperation of a mobile robot or using a hand-held camera, such as video sequences from a RGB-D camera. Therefore, these systems are not fully autonomous and depend on human control to explore and map environments.
Generally speaking, these methods fuse semantic information about objects into a 3D map from an SLAM method. For example, SemanticFusion [13] combines ElasticFusion [31] for dense reconstruction with a neural network for semantic segmentation. DA-RNN [24] combines KinectFusion [9] and a recurrent neural network for video semantic segmentation. [12] uses Mask R-CNN [32] for instance segmentation to construct and update a semantic map in real-time using a mobile manipulator system. However, it requires a fully explored environment map as a prerequisite to begin with. Hydra [25] builds and optimizes a 3D scene graph hierarchically from simulated and real-world datasets. Khronos [26] constructs dense spatio-temporal maps that reflect the 3D scene and also account for some dynamic environment changes while testing their method using a physical robot system. However, robots are manually controlled during this process. Recently, open-vocabulary object detectors such as GroundingDINO [33] have been utilized for semantic mapping. For example, ConceptGraphs [14] and its hierarchical version [34] build 3D scene graphs with open-vocabulary object detectors based on a hand-h eld scanned dataset of the indoor environment, having ground truth camera poses.
Although these works significantly advance in semantic mapping, they all share a common limitation, i.e., they are not fully autonomous. As shown in Table I, prior methods are based on manual control for environmental exploration. Some methods require pre-explored maps, and some do not perform semantic updating of environments. In contrast, our method initially builds the map through robot exploration.
Figure 2: Illustration of our system for autonomous exploration, semantic mapping and map update.
Consequently, the robot can use this map to revisit the environment in a planned and greedy manner, while continuously refining and updating the semantic information in the environment. This offers a fully autonomous approach to both mapping and semantic updating.
III System
Our system consists of two phases. In the initial phase, a mobile robot explores the environment and builds the occupancy map incrementally. In the second phase, the robot leverages the built occupancy map to accurately localize within the environment and performs semantic mapping. The motivation to separate the process into two stages, rather than performing simultaneously, is because of computational efficiency and robustness. Performing both tasks in a single stage leads to resource constraints and reduced performance, especially in large-scale environments with many objects. Additionally, in large environments, as the number of detected objects increases, the number of map corrections (like loop closures) applied to the detected objects increase, which slows down the mapping process. This has also been mentioned in [26]. Therefore, our two-phase approach is more suitable for semantic mapping of large-scale environments. Furthermore, this separation provides a significant advantage for Lidar-based mapping that relies on scan matching, such as GMapping [35] and Hector SLAM [36]. Robot poses are not reliably maintained during the initial mapping for these methods as they rely on scan-matching. Accurate robot poses are obtained during the localization within a built map. Our two-stage approach allows these techniques to benefit from a reliable robot pose and thus more reliable object poses during the semantic mapping stage. An overview of our system is presented in Fig. 2.
III-A Autonomous Exploration and Map Building
We have developed an environment exploration method based on the popular frontier exploration module [37], using a dynamic search window to efficiently navigate to frontiers in large real-world environments. When using the original approach in large-scale environments, as the map size increases, far-away frontiers become insignificant due to high cost, thus leaving a part of the map unexplored. Moreover, tuning the cost parameters in the original method is not intuitive for large environments.
Dynamic Window Frontier Exploration (DWFE). We modified the original frontier exploration method  [37] by dynamically adjusting the robot’s search area between local and global regions, as detailed in Algorithm 1222Code available at https://github.com/IRVLUTD/dynamic-window-frontier-exploration. This method identifies frontiers in the current search area and the robot navigates towards them to progressively map the environment. The search window is adjusted dynamically based on the density of the frontiers available around the robot. If the number of frontiers within the local search radius around the robot falls below a threshold, the search expands to a global radius, as seen in Fig. 3. Setting this parameter to a large value expands the search region to the entire map. This allows the robot to move progressively ahead, covering the local regions well, reducing redundant coverage of the environment. The exploration continues until the number of frontiers falls below a global threshold or after a fixed time has elapsed, at which point the exploration terminates and saves the built map. Our improved exploration algorithm is integrated with the GMapping SLAM module [35, 38]. The approach is depicted in Fig. 2(a). During this process, the robot exploration trajectory is recorded, which is later used to obtain a traversal plan. The saved occupancy map is then utilized to localize the robot during the semantic map construction and update phases.
Algorithm 1 Dynamic Window Frontier Exploration
1:Input: Occupancy map $M$, local, global radii $r_{l}$, $r_{g}$, local, global threshold $T_{l}$, $T_{g}$, max time $T_{max}$, weights $\alpha,\beta$
2:Initialize robot position $p$, time $t\leftarrow 0$
3:while $t<T_{max}$ do
4:  $F\leftarrow$ frontiers within $r_{l}$ around $p$
5:  if $|F|<T_{l}$ then
6:   $F\leftarrow$ frontiers within $r_{g}$ around $p$
7:   if $|F|<T_{g}$ then
8:     break $\triangleright$ Terminate if not enough frontiers
9:   end if
10:  end if
11:  $f^{*}\leftarrow\arg\min_{f\in F}\left(\alpha\cdot distance(p,f)+\beta\cdot\text{size}(f)\right)$
12:  yield $f^{*}$ $\triangleright$ Output the best frontier for this iteration
13:  Navigate to $f^{*}$
14:  $t\leftarrow t+1$
15:end while
Figure 3: Our dynamic search window during exploration.
Figure 4: Environment traversal trajectory planning.
III-B Autonomous Semantic Mapping
We utilize the saved occupancy map and the recorded robot exploration path to plan a trajectory to traverse the environment again swiftly, to construct the semantic map.
III-B1 Environment Traversing
To construct or update a semantic map, the robot needs to visit places in the environment and identify the objects. To optimize this travel, we first sample a subset of $n$ points $P=\{p_{1},p_{2},\dots,p_{n}\}$ along the robot exploration path (Fig. 4(a)), with a minimum of $2m$ separation between each pair of points (Fig. 4(b)). We then construct a graph $G(P,E)$ with the sampled way points as nodes and the Euclidean distance $d$ between them as corresponding edge weights $w_{ij}=d_{ij}=\|p_{i}-p_{j}\|$. This is depicted in Fig. 4(c). Finally, a traversal trajectory $T_{s}$ is obtained from $P$ using a greedy formulation of the Traveling Salesmen Problem (TSP) as explained in Algorithm 2. The generated trajectory is shown in Fig. 4(d).
Algorithm 2 Greedy Algorithm for TSP
1:Input: Sampled waypoints $P=\{p_{1},p_{2},\dots,p_{n}\}$, where $p_{1}$ is the starting waypoint
2:Output: A sequence of waypoints representing the TSP solution trajectory $T_{s}$
3:Initialize trajectory $T_{s}$ $\leftarrow[p_{1}]$
4:Initialize visited $\leftarrow\{p_{1}\}$
5:for $t\leftarrow 1$ to n1n-1 do
6:  Let $p_{\text{last}}\leftarrow$ last waypoint in the $T_{s}$
7:  $p_{\text{next}}\leftarrow\arg\min_{p_{j}\in P\setminus\text{visited}}w(p_{\text{last}},p_{j})$
8:  Append $p_{\text{next}}$ to $T_{s}$
9:  Add $p_{\text{next}}$ to visited
10:end for
11:Append $p_{1}$ to $T_{s}\triangleright$ Return to the starting waypoint
12:Return trajectory $T_{s}$
III-B2 Object Detection and Segmentation
During traversal, object detection and segmentation are performed in real time to build the semantic map. Our system utilizes GroundingDINO [33] to detect objects from the robot’s RGB image observation, providing labels and bounding boxes, which are then used as prompts for MobileSAM [39], a faster version of SAM [40], to generate segmentation masks.
After obtaining object segments, we compute the 3D point cloud $\mathbf{P}_{\text{camera}}$ for each segment in the camera frame by applying the intrinsic parameters of the camera to the depth image observation. These points are then transformed into the map frame $M$, through the transformation matrices representing the camera pose in the robot base frame, $\mathbf{T}_{\text{camera}}\in\mathbb{SE}(3)$ and the robot pose in the map frame $\mathbf{T}_{\text{robot}}\in\mathbb{SE}(3)$, as follows: $\mathbf{P}_{\text{map}}=\mathbf{T}_{\text{robot}}\cdot\mathbf{T}_{\text{camera}}\cdot\mathbf{P}_{\text{camera}}$. Finally, the mean position of the point cloud is computed and represented as the 3D position of the corresponding object in the map frame.
III-B3 Semantics Representation and Mapping
Our semantic map is a hybrid representation. At the lower level, we utilize a 2D occupancy grid map $M$ that captures the spatial information of the free space and obstacles in the environment. At the higher level, we construct a topological map defined by a set of vertices $V=\{v_{1},v_{2},\dots,v_{n}\}$. Each vertex $v_{i}$ represents a uniquely detected objects (Fig. 2(d)). Each node $v_{i}$ has the following attributes.
  • $v_{i}.\text{id}$: A unique id of the object
  • $v_{i}.\text{category}$: Object category label
  • $v_{i}.\text{position}$: Mean position of the object computed in the reference frame of the occupancy map. This attribute connects the topological map to the occupancy map in a hierarchical way.
  • $v_{i}.\text{confidence}$: the confidence score of the detected object from object detection
This representation facilitates fast object association and easy update of nodes during environment changes.
Semantic Mapping and Updating. To build the semantic map, we first load the occupancy map $M$ built as detailed in Sec. III-A and localize the robot within it. If the semantic mapping process is performed for the first time, a topological map is initialized with an empty set of vertex $V^{\prime}=\emptyset$. Otherwise, the previously constructed semantic map is loaded ( $V^{\prime}\leftarrow V$). For each object category (e.g., tables, chairs, doors), a category-specific object association distance threshold $\delta_{\{\text{category}\}}$ is defined to determine whether a detected object corresponds to an existing one. Categories without explicit thresholds are assigned a default value. In practice, setting $\delta$ roughly equal to the typical object size improves the differentiation between nearby objects.
Algorithm 3 Semantic Mapping and Updating
1:Input: Traversal trajectory $T_{s}$, Distance thresholds $\Delta=\{\delta_{c}\mid c\in\text{categories}\}$
2:Output: Updated semantic map $G_{T}(V^{\prime})$
3:if no prior map exists then
4:  Initialize $V^{\prime}\leftarrow\emptyset$, id $\leftarrow 0$
5:else
6:  Load $G_{T}(V)$
7:  $V^{\prime}\leftarrow V$,  $\text{id}\leftarrow\max_{v\in V^{\prime}}v.\text{id}+1$
8:end if
9:while Traversing $T_{s}$ do
10:  $\mathcal{O}_{\text{current}}\leftarrow$ objects detected in FoV
11:  $\mathcal{V}_{\text{expected}}\leftarrow$ expected nodes in FoV from $G_{T}(V^{\prime})$
12:  $\mathcal{O}_{\text{unmatched}}\leftarrow\mathcal{O}_{\text{current}}$ $\triangleright$ Initialize unmatched objects
13:  for $v\in\mathcal{V}_{\text{expected}}$ do
14:   if no $o\in\mathcal{O}_{\text{current}}$ such that
[…]
17:     $V^{\prime}\leftarrow V^{\prime}\setminus\{v\}$ $\triangleright$ Remove node
18:   else
19:     $\mathcal{O}_{\text{unmatched}}\leftarrow\mathcal{O}_{\text{unmatched}}\setminus\{o\}$
20:   end if
21:  end for
22:  for $o\in\mathcal{O}_{\text{unmatched}}$ do
23:   $V_{c}\leftarrow\{v\in V^{\prime}\mid v.\text{category}=o.\text{category}\}$
24:   $d_{\min}\leftarrow\infty$
25:   if $V_{c}\neq\emptyset$ then
26:     $d_{\min}\leftarrow\min_{v\in V_{c}}\text{dist}(P_{\text{mean}}(o),v.\text{position})$
27:   end if
28:   if $d_{\min}>\delta_{o.\text{category}}$ then
29:     $v_{new}=(\text{id},o.\text{category},$
30:          $P_{\text{mean}}(o.\text{segment}),$
31:          $o.\text{confidence})$
32:     $V^{\prime}\leftarrow V^{\prime}\cup\{v_{new}\}$ $\triangleright$ Add new node to $V^{\prime}$
33:     $\text{id}\leftarrow\text{id}+1$
34:   end if
35:  end for
36:end while
37:Return $G_{T}(V^{\prime})$
Once initialized, the robot then begins following the traversal trajectory $T_{s}$ (see Sect. III-B1). Simultaneously, the object detection and segmentation module processes the current image observation of the robot’s field of view (FoV) to detect objects $\mathcal{O}_{\text{current}}$ and compute their positions in the map $M$ frame of reference as seen in Fig. 5(a). These detections are initially considered unmatched $\mathcal{O}_{\text{unmatched}}$.
At each step, an approximate polygon representing the field of view is derived from the limits of the current depth point cloud. Using this FoV polygon, the semantic map is queried to retrieve the nodes $\mathcal{V}_{\text{expected}}$, which represent the objects expected to be within the FoV. The Semantic Mapping procedure (described in Alg. 3 iteratively updates the vertex set $V^{\prime}$.
Node Removal. For each expected node $v\in\mathcal{V}_{\text{expected}}$, the algorithm checks for a matching object $o\in\mathcal{O}_{\text{current}}$ of the same category ($o.\text{category}=v.\text{category}$) and within the threshold distance $\text{dist}(P_{\text{mean}}(o),v.\text{position})\leq\delta_{v.\text{category}}$. If no match is found, the node $v$ is removed from $V^{\prime}$, indicating that the corresponding object is no longer present in the expected location (e.g., moved or removed). The matched objects are removed from $\mathcal{O}_{\text{unmatched}}$.
Node Addition. For each unmatched object $o\in\mathcal{O}_{\text{unmatched}}$, the algorithm identifies nodes $V_{c}=\{v\in V^{\prime}\mid v.\text{category}=o.\text{category}\}$ belonging to the same category. If no such category exists $V_{c}=\emptyset$ or an object is not associated with existing ones $\min_{v\in V_{c}}\text{dist}(P_{\text{mean}}(o),v.\text{position})>\delta_{o.\text{category}}$, a new node $v_{new}=(\text{id},o.\text{category},P_{\text{mean}}(o.\text{segment}),o.\text{confidence})$ is added to $V^{\prime}$. The object association mechanism is illustrated in Fig. 5(b). This iterative process continues until the robot completes the traversal of $T_{s}$. By continuously updating the vertex set $V^{\prime}$, the algorithm ensures that the semantic map captures both short-term and long-term environmental changes. The resulting map seamlessly integrates the geometric structure of the occupancy map with the semantic information of the topological map , enabling robust and faster scene understanding.
Figure 5: Updating Semantic Map through Object Association.
IV Experiments
To evaluate the effectiveness of our proposed Dynamic Window Frontier Exploration method (Alg. 1), we conducted experiments in a large-scale simulated hospital environment[41] of approximately $1440~\text{sq.m}$, in the Gazebo simulator. We conducted 20 iterations of exploration, where in each iteration a Fetch mobile manipulator robot was spawned at a different starting location and explored for 15 minutes. The same starting locations are used for both our approach and conventional frontier exploration. During each run, we recorded the area covered and the path length traveled, allowing a direct comparison between our approach and the conventional frontier exploration baseline [37].
[…]
Baseline[37] 669.12 107.10
DFWE (ours) 721.05 94.30
TABLE II: Average area coverage and path length for baseline vs. DFWE in a large-scale simulated environment.
The results in Table II demonstrate the improved exploration efficiency using our method. Compared to the baseline exploration, our method explores $\mathbf{\approx 7.8\%}$ more area while traveling $\mathbf{\approx 12\%}$ less on average. These numbers are expected to get better as the size of environment increases.
We conducted semantic mapping experiments in two different real-world environments with different object categories and evaluated the performance of our system.
Environment A. This is a large indoor environment measuring approximately $\mathbf{93m\times 90m}$. It consists of 18 corridors with a total traversal length of approximately 800 m These corridors were mostly furnished with typical office furniture, including doors, tables, and chairs. Therefore, we only consider these 3 classes, i.e., table, door and chair for reliable semantic mapping of this environment.
Environment B. This is a laboratory environment with a reduced size measuring $9m\times 13m$. This space included a more diverse set of objects such as televisions, trash bins, umbrellas, chairs, cabinets and persons. Only objects with a detection confidence of 0.6 or higher are considered.
Robot. We run the experiments on a Fetch mobile robot equipped with a LiDAR sensor and an RGB-D camera.
IV-A Real World Exploration and Mapping
We integrate mapping, exploration, and navigation to autonomously map environments. Using Gmapping [38], a 0.1 m/pixel resolution occupancy map was built using laser scans, while our Dynamic Window Frontier Exploration module selected the best frontiers in the search radius. The move_base ROS package [42] guided the robot safely to these frontiers with global and local planning, dynamically avoiding unexpected obstacles. The robot was limited to 0.6 m/s speed with a 0.7 m inflation radius for safe navigation in the environment. In Environment A, 150 min of the exploration (including stops for people in narrow corridors) produced a full map and 2,250 trajectory points in 0.25 Hz. In Environment B, the exploration was completed in approximately 4 min, recording 212 points at 1 Hz.
Environment Traversing. Using the greedy TSP from Sec. III-B1, traversal trajectories were generated from the recorded exploration poses: 130 poses for Environment A and 15 poses for Environment B. Revisiting was significantly faster, taking approximately 35 min for Environment A and 2 min for Environment B, demonstrating the efficiency of the optimized traversal.
IV-B Real-time Semantic Mapping
For real-time semantic mapping, we equip Fetch with a laptop with an RTX 4090 GPU running Ubuntu 20.04 and ROS Noetic. The system then begins with the built occupancy map. After localizing using AMCL ROS [42], the robot progressively builds the semantic map. We set the object association distance threshold of 0.7 m for all categories. Subsequently, two additional iterations are performed, during which the environment is modified through the removal, addition, and/or relocating objects in the environments. Fig. 6 shows some of the changes in Environment B, reflected during semantic updating.
Env Class Iteration I (Initial run) Iteration II (vs I) Iteration III (vs I)
   GT Map $\DeltaGT Δ\DeltaMap$ $\DeltaGT Δ\DeltaMap$
A Table 45 45 0 -6 0* -8
Chair 109 84 -10 -11 0* -15
Door 170 153 0 +2 0 +3
B Person 2 2 0* 0 0 0
Trash Bin 2 2 -2 -2 -2 -1
Umbrella 1 1 -1 -1 -1 -1
Chair 3 3 0 0 -2 -2
Television 1 2 0 -1 0 -1
Cabinet 3 2 0 -1 0 0
[…]
Figure 6: Visualization of environmental changes in Environment B: Iteration II (2 trash bins removed) and Iteration III (2 chairs removed). When robot revisited, these changes are reflected in the updated semantic map.
IV-C Quantitative Analysis.
Table III reports the semantic mapping results across three iterations, with Iteration I serving as the reference. For Iteration I, we show the absolute ground truth (GT) objects and those mapped by the robot (Map). For Iterations II and III, we report only the relative changes ($\DeltaGT, Δ\DeltaMap$) compared to Iteration I.
[…]
Overall, this analysis shows that while stable objects (e.g., tables, doors) maintain consistent mapping across iterations, dynamic or occlusion-prone categories such as chairs exhibit greater variability in detection and mapping performance.
Class Iteration I Iteration II Iteration III
TP FP FN F1 TP FP FN F1 TP FP FN F1
Environment A
Table 36 9 9 0.8 36 3 9 0.86 33 4 12 0.80
Chair 84 0 25 0.87 72 1 27 0.84 67 2 42 0.75
Door 140 13 30 0.87 140 15 30 0.86 142 14 28 0.87
Environment B
Person 2 0 0 1.00 2 0 0 1.00 2 0 0 1.00
Trash Bin 2 0 0 1.00 0 0 0 - 0 1 0 0.00
Umbrella 1 0 0 1.00 0 0 0 - 0 0 0 -
Chair 3 0 0 1.00 3 0 0 1.00 1 0 0 1.00
Television 1 1 0 0.67 1 0 0 1.00 1 0 0 1.00
Cabinet 2 0 1 0.80 1 0 2 0.50 2 0 1 0.80
TABLE IV: Semantic map object count across Iteration I, II & III for A & B Environments. TP: True Positives, FP: False Positives, FN: False Negatives, F1 scores are reported.
Figure 7: Intermediate semantic maps generated by : (a) our method, (b) Khronos [26], running at 21FPS of data replay.
We compare and evaluate our semantic mapping approach with Khronos [26] on the real-world mezzanine scene dataset [43] provided as a ROS bag. Fig. 7 represents the semantic maps generated by both methods at an intermediate stage. The results are summarized in Table V. At 30 FPS of ROS bag replay, our method successfully maps 79 objects, while Khronos only maps 3, showing that our approach can process visual information faster and efficient as well. When the input rate is reduced to 21 FPS, Khronos maps 87 objects, while our method achieves a comparable 86.
Data Replay Rate Objects Mapped by Method
Khronos [26] Ours
21 FPS 87 86
30 FPS 3 79
TABLE V: Number of objects mapped by each method at different data FPS on the real-world mezzanine dataset.
This performance gap arises from the fundamental differences in system design. Since Khronos performs SLAM and semantic mapping in a single stage, its runtime is to an extent governed by loop-closure optimization, achieving superior SLAM performance, but limits the scalability of semantic mapping at higher frame rates. In contrast, our method decouples occupancy and semantic mapping, allowing it to achieve both higher throughput and competitive accuracy across different operating conditions.
IV-D Qualitative Analysis
Figure 8: Examples of accurate detection and segmentation of chairs, tables and doors in Environment A.
Fig. 8 illustrates several examples of accurate object detection during our experiments for chairs, tables, and doors.
In Fig. 9, we show several challenging scenarios in object detection and segmentation, while robot is navigating.
Occlusion Challenges. Close proximity of objects such as tables and chairs causes occlusion and leads to missed detections or incorrect segmentation, as seen in Fig. 9(a). In Fig. 9(c) table, chair, and a portion of the pillar are segmented together as table. The scenario seen in Fig. 9(b), where a chair is partially occluded by the pillar, leads to incorrect object associations when the chair is detected again at another view.
Detection Threshold. Lowering the detection threshold increases false positives, while a higher threshold may miss objects. In Fig. 9(e), chairs in the foreground are undetected at a 0.8 threshold. careful choice of threshold for object class is important for accurate semantic mapping.
Limitations in Constrained Spaces. Narrow corridors limit the robot’s view of objects like doors, as seen in Fig. 9(f). Several of such doors are not detected, leading to a reduction in the number of object nodes in the semantic map. This is evident in Table III.
Figure 9: Representation of various challenging scenarios involved in object detection and semantic mapping.
Deviation from the Traversal Trajectory. In Fig. 9(d), a large bin in the corridor occluded objects, forcing the robot to take a detour. This detour resulted in the robot not seeing some objects. Such scenarios impact the quality of the semantic mapping, reducing the ability to accommodate environmental changes.
Impact of Lighting. Varying lighting conditions can cause shadows and reflections. As seen in Fig. 9(g) and Fig. 9(h), these shadows on the floor were mistook for a door and a table, respectively, reducing the overall detection and semantic map accuracy. In Fig. 9(c), reduced brightness at night degraded detection and segmentation.
Structural Misidentification. Walls resembling doors due to shape or added features (e.g., support bars) often cause misidentification, as seen in Fig. 9(i). This scenario also occurred in Environment B, update II phase, where a cardboard box was identified as a trash bin, leading to a false positive.
V Conclusion and Future Work
We introduce a robotic system that enables a mobile robot to autonomously explore an unknown environment, recognize objects in the environment, and subsequently build a semantic map of the environment. The system utilizes a hybrid representation of the semantic map that consists of an occupancy grid map for geometry and a topological map for semantics. This representation enables us to update the semantic map efficiently to reflect the changes in the environment. Experiments conducted on a mobile manipulatory system in large- and medium-scale indoor environments demonstrate the effectiveness of our system in autonomous exploration and semantic mapping.
Object perception during robot navigation still remains a challenge. Our future work will consider interactive object recognition, in which a robot can plan its actions to detect objects such as looking at some objects, moving towards an object to confirm its existence for a robust object association, and dynamic planning of traversal to avoid occlusions.
Acknowledgment
This work was supported by the DARPA Perceptually-enabled Task Guidance (PTG) Program under contract number HR00112220005, the Sony Research Award Program, the National Science Foundation (NSF) under Grant Nos. 2346528 and 2520553, and the NVIDIA Academic Grant Program Award. The work of T. Summers was supported by the United States Air Force Office of Scientific Research under Grant FA9550-23-1-0424 and the National Science Foundation under Grant ECCS-2047040. We would like to thank our colleague, Jishnu Jaykumar P, for his assistance during the experiments.